# EVA-MILP: Towards Standardized Evaluation of MILP Instance Generation

### v2

@@ Appendix A Datasets
Ecole Synthesized Instances
To facilitate various experiments, including model training and evaluation across different metrics, 4 synthetic datasets of MILP instances was generated. This process utilized the Ecole library [28], a platform designed for machine learning research in combinatorial optimization.
  • Configuration File: Instance generation is controlled via a YAML configuration file. This file defines:
    • Global Settings: A random seed for reproducibility, the base output directory, and the total number of instances to generate per problem type.
    • Size Ranges: Default minimum and maximum values for the number of constraints, number of variables, and density, applicable across all problem types unless overridden.
    • Problem-Specific Settings: Sections dedicated to each problem type allow for enabling/disabling generation and overriding global size ranges. They also accommodate problem-specific parameters, such as the ratio of facilities to customers in CFL or n_items in CA.
  • Parameter Randomization: Parameters like the number of constraints, variables, and density are randomly sampled for each instance within the ranges specified in the configuration file. This ensures diversity in the generated dataset. The configuration allows setting both global ranges and more specific ranges per problem type. The details of synthesizing parameters are in 6.
Compatibility Testing with Public Datasets (NIPS ML4CO)
In addition to utilizing the synthetically generated dataset described above, the compatibility and robustness of the various analysis pipelines and tools developed (e.g., for feature extraction, solver log parsing, model evaluation) were verified using publicly available MILP benchmark instances. Specifically, datasets from the ML4CO Competition [29] were employed for testing purposes. Successfully processing and running analyses on these standard benchmarks demonstrated the broader applicability and compatibility of the developed experimental framework beyond the specific synthetically generated instances.
@@
Appendix B Related Work
@@
  • VAE-based Approaches: G2MILP [10] introduced the first DL-based framework, using a masked Variational Autoencoder (VAE) paradigm. It iteratively corrupts and replaces constraint nodes in the bipartite graph representation. While capable of generating instances structurally similar to the training data, G2MILP does not explicitly guarantee the feasibility or boundedness of the generated instances, potentially leading to unusable samples for certain downstream tasks. DIG-MILP [11], also VAE-based, addresses the feasibility and boundedness guarantee by leveraging MILP duality theories and sampling from a space including feasible solutions for both primal and dual formats. However, this can come at the cost of structural similarity compared to the original data. ACM-MILP [12] refines the VAE approach by incorporating adaptive constraint modification. It uses probability estimation in the latent space to select instance-specific constraints for modification, preserving core characteristics. It also groups strongly related constraints via community detection for collective modification, aiming to maintain constraint interrelations and improve hardness preservation.
  • Diffusion-based Approaches: Recognizing the power of diffusion models in generation tasks, MILP-FBGen [16] proposed a diffusion-based framework. It uses a structure-preserving module to modify constraint/variable vectors and a feasibility/boundedness-constrained sampling module for the right-hand side and objective coefficients, aiming to ensure both structural similarity and feasibility/boundedness. Another diffusion-based approach focuses specifically on generating feasible solutions (rather than full instances) using guided diffusion. It employs contrastive learning to embed instances and solutions and uses a diffusion model conditioned on instance embeddings to generate solution embeddings, guided by constraints and objectives during sampling.
  • Structure-Focused Approaches: Some recent works explicitly target specific structural properties. MILP-StuDio [13] focuses on preserving and manipulating block structures commonly found in CCMs of real-world MILPs. It decomposes instances into block units, builds a library of these units, and uses operators (reduction, mix-up, expansion) to construct new, potentially scalable instances while maintaining structural integrity and properties like hardness and feasibility. MILPGen [41]also aims at scalability, using node splitting/merging on the bipartite graph representation to decompose instances into tree structures, which are then concatenated to build larger problems.
@@ Appendix C Parameter Summary
See the parameters of problems in Table 7, 8, 9.
Table 7: Summary of Key Experimental Parameters (Part 1 of 3).
@@
Feasibility Ratio
Gurobi time_limit 300 s Maximum Gurobi solve time per instance for feasibility/boundedness check.
Gurobi threads 64 Number of CPU threads used per Gurobi solve.
parallel true Flag enabling parallel processing of instances.
n_jobs null Number of parallel workers (default: CPU cores - 1).
Structural Similarity
Feature Set 11 features Set of graph-based structural features extracted from each MILP instance (e.g., density, degree stats, clustering).
num_workers 64 Number of parallel processes for feature extraction.
num_samples 1000 Number of instances sampled (with replacement) from each dataset for JSD calculation.
Comparison Metric JSD-based Jensen-Shannon Divergence between feature distributions, converted to similarity ($S_{i}=1-JSD_{i}/\log(2)$) and averaged.
Branching Nodes (Hardness Comparison)
Gurobi time_limit 300 s Maximum Gurobi solve time per instance.
Gurobi threads 1 Threads per Gurobi solve (to enhance determinism for node count comparison).
Gurobi Heuristics 0.5 Gurobi heuristic effort parameter setting.
Gurobi MIPFocus 1 Gurobi parameter to focus on finding feasible solutions.
Metric Relative Error Relative difference in total B&B nodes: $|(\sum N_{gen}-\sum N_{train})/\sum N_{train}|\times 100\%$.
Solving Time Gap (Hardness Comparison)
Gurobi time_limit 100 s Maximum Gurobi solve time per instance.
Gurobi threads 256 Threads per Gurobi solve (high count used to expedite experiments).
[…]
Table 8: Summary of Key Experimental Parameters (Part 2 of 3).
Category Parameter Name Value Description
Hyperparameter Tuning (SMAC3 for Gurobi)
Tuning Framework SMAC3 Automated algorithm configuration tool used.
n_trials 50 SMAC3 evaluation budget (number of configurations tested).
random_seed 42 Seed for SMAC reproducibility.
Objective Mean Solve Time Minimize mean wall-clock time on the tuning instance set.
Gurobi time_limit 60 s Gurobi time limit per solve during tuning and final evaluation.
Gurobi threads 256 Gurobi threads per solve during tuning and final evaluation.
Tuned Parameters (8) Gurobi Controls Heuristics, MIPFocus, VarBranch, BranchDir, Presolve, PrePasses, Cuts, Method. (Ranges/values specified in config).
Initial Basis Prediction (GNN)
GNN Model Parameters
@@
epochs 800 (Max) Maximum number of training epochs.
batch_size 32 Number of instances per training batch.
learning_rate 1e-3 Initial learning rate for the Adam optimizer.
weight_decay 1e-4 L2 regularization coefficient.
early_stopping 50 Patience (epochs) for early stopping.
seed 42 Random seed for training reproducibility.
@@
Gurobi threads 256 Gurobi threads per solve during evaluation.
Gurobi mip_gap 0.01 Gurobi target MIP gap during evaluation.
@@
Basis Repair Iterations 50 Max iterations for basis repair algorithm.
Table 9: Summary of Key Experimental Parameters (Part 3 of 3).
Category Parameter Name Value Description
Solver-Internal Features
Data Generation Phase (Gurobi Run)
Gurobi time_limit 100 s Gurobi time limit per instance for log generation.
Gurobi threads 128 Gurobi threads per instance for log generation.
Gurobi seed 42 Gurobi seed for log generation reproducibility.
Gurobi mip_gap 1e-4 Gurobi MIP gap for log generation runs.
Callback Usage Enabled Gurobi callback potentially used for detailed heuristic data.
@@
Root Gap / Heuristic Metric 1-Wasserstein Dist. Metric used to compare scalar distributions.
Cut Plane Metric PCA + 1-Wasserstein Method for comparing high-dim cut vectors.
@@
PCA Scaling Standard Data scaling method (Z-score) applied before PCA.
@@ Appendix E More Preliminaries
E.1 Fundamental Optimization Concepts
Feasibility  Consider the MILP problem defined by constraints $\mathbf{A}\mathbf{x}\leq\mathbf{b}$ and $x_{i}\in\mathbb{Z},\forall i\in\mathcal{I}$. The feasible region is the set =$\mathcal{F}=\{\mathbf{x}\in\mathbb{R}^{n}\mid\mathbf{A}\mathbf{x}\leq\mathbf{b},x_{i}\in\mathbb{Z}\;\forall i\in\mathcal{I}\}$. The MILP problem is feasible if its feasible region $\mathcal{F}$ is non-empty, i.e., $\mathcal{F}\neq\emptyset$. A specific point $\hat{\mathbf{x}}\in\mathbb{R}^{n}$ is considered feasible if it satisfies all constraints, i.e., $\hat{\mathbf{x}}\in\mathcal{F}$.
E.2 Core Solution Algorithms and Techniques
@@
Simplex Method Basics  The Simplex Method [42] is an algorithm primarily designed for solving Linear Programming (LP) problems, such as the LP relaxation defined above. While it can handle various forms, it often operates by converting the problem to a standard form like $\min\{\mathbf{c^{\prime}}^{\top}\mathbf{y}\mid\mathbf{A^{\prime}}\mathbf{y}=\mathbf{b^{\prime}},\mathbf{y}\geq 0\}$ (where $\mathbf{y}$ might include original and slack/surplus variables). It explores the vertices (Basic Feasible Solutions, BFS) of the feasible polyhedron $\mathcal{P}^{\prime}=\{\mathbf{y}\mid\mathbf{A^{\prime}}\mathbf{y}=\mathbf{b^{\prime}},\mathbf{y}\geq 0\}$. It iteratively moves between adjacent BFS by exchanging one variable in the current basis (a set of variables determining the BFS) with a non-basic variable, guided by criteria like reduced costs to ensure monotonic improvement of the objective function value, until an optimal BFS is found or unboundedness is detected.
Heuristics  Within MILP solvers, heuristics [35] are algorithms designed to rapidly find a feasible integer solution $\hat{\mathbf{x}}\in\mathcal{F}_{MILP}=\{\mathbf{x}\mid\mathbf{A}\mathbf{x}\leq\mathbf{b},x_{i}\in\mathbb{Z}\;\forall i\in\mathcal{I}\}$, whose objective value $\mathbf{\tilde{c}}^{\top}\hat{\mathbf{x}}$ is hopefully close to the true optimum $z_{MILP}$. They do not guarantee optimality. Their main purpose is to quickly establish strong primal bounds (updating the incumbent $z_{incumbent}=\min(z_{incumbent},\mathbf{\tilde{c}}^{\top}\hat{\mathbf{x}})$). In branch-and-bound, these incumbents enable pruning of search nodes $j$ where the local lower bound $z_{bound}^{(j)}$ satisfies $z_{bound}^{(j)}\geq z_{incumbent}$.
Cutting Planes (Cuts)Cutting planes [43] are linear inequalities, $\alpha^{T}\mathbf{x}\leq\beta$, that are valid for all feasible integer solutions, i.e., they hold for all $\mathbf{x}\in Conv(\mathcal{F}_{MILP})$ where $\mathcal{F}_{MILP}=\{\mathbf{x}\mid\mathbf{A}\mathbf{x}\leq\mathbf{b},x_{i}\in\mathbb{Z}\;\forall i\in\mathcal{I}\}$. However, they are chosen such that they are violated by the optimal solution $\mathbf{x}^{*}_{LP}$ of the current LP relaxation (i.e., $\alpha^{T}\mathbf{x}^{*}_{LP}>\beta$). Adding such cuts $\mathcal{C}$ to the LP relaxation feasible set $\mathcal{P}=\{\mathbf{x}\mid\mathbf{A}\mathbf{x}\leq\mathbf{b}\}$ yields a tighter relaxation $\mathcal{P}^{\prime}=\mathcal{P}\cap\{\mathbf{x}\mid\alpha^{T}\mathbf{x}\leq\beta$ for all, $\alpha^{T}\mathbf{x}\leq\beta\in\mathcal{C}\}$. The goal is to improve the lower bound derived from the relaxation, $\min_{\mathbf{x}\in\mathcal{P}^{\prime}}\mathbf{\tilde{c}}^{\top}\mathbf{x}\geq z_{LP}$, thereby bringing it closer to the true MILP optimum $z_{MILP}$ and potentially reducing the search space.
Appendix F Feasibility Ratio
The primary goal of this experiment is to calculate the Feasibility Ratio for a given dataset of Mixed-Integer Linear Programming (MILP) instances. The Feasibility Ratio is defined as the percentage of instances within the dataset that are determined to be both feasible (possessing at least one valid solution satisfying all constraints) and bounded (having a finite optimal objective value, not tending towards infinity or negative infinity). Instances that are infeasible, unbounded, or whose status cannot be determined within the given time limit are not counted towards this ratio.
Methodology & Operations
  1. 1.
    Key parameters loaded include:
    • time_limit: 300 seconds (Maximum solving time per instance).
    • threads: 64 (Number of CPU threads Gurobi can use per solve).
  2. 2.
    Feasibility and Boundedness Check: The instance is considered Feasible and Bounded for the metric only if the status is GRB.OPTIMAL (2) or GRB.SUBOPTIMAL (12). It is considered not meeting the criteria if the status is GRB.INFEASIBLE (3), GRB.INF_OR_UNBD (4), GRB.UNBOUNDED (5), or GRB.TIME_LIMIT (9). Any other status or error during the process also results in the instance being considered not feasible/bounded.
  3. 3.
    Result Aggregation:
    • After all instances have been checked (either sequentially or in parallel), the results (a boolean indicating feasible/bounded status for each instance) are collected.
    • The total number of instances processed (total_count) and the number of instances deemed feasible and bounded (feasible_count) are tallied.
    • The Feasibility Ratio is calculated as:
      Feasibility $\text{Feasibility Ratio}=\left(\frac{\texttt{feasible\_count}}{\texttt{total\_count}}\right)\times 100$
  4. 4.
    Output Generation:
    • A results report file is automatically generated within this directory.
    • This report contains:
      • The input instance directory path.
      • The total number of instances found.
      • The number of instances found to be feasible and bounded.
      • The calculated Feasibility Ratio (%).
      • A detailed list of each instance file name and its corresponding status ("Feasible and bounded" or "Infeasible or unbounded").
    • A summary of the results (directory, counts, ratio, and output file path) is also printed to the console upon completion.
Summary of Parameters
  • Solver Time Limit per Instance: 300 seconds
  • Solver Threads per Instance: 64
  • Execution Mode: Parallel
  • Number of Parallel Workers: System CPU core count minus 1
  • Metric Definition: Percentage of instances with Gurobi status OPTIMAL or SUBOPTIMAL within the time limit.
  • Output: Text report and console summary.
Table 10: Feasible and Bounded Ratio for IS Instance Sets
Dataset / Model Masked Ratio Total Instances Feasible & Bounded
Ratio (%)
Original Training Set (raw) N/A 1000 100.00%
G2MILP (g2milp_mis) 0.01 1000 100.00%
0.05 1000 100.00%
0.1 1000 93.40%
ACM-MILP (acmmilp_mis) 0.05 1000 100.00%
0.1 1000 100.00%
0.2 1000 100.00%
Table 11: Feasible and Bounded Ratio for Combinatorial Auction (CA) Instance Sets
Dataset / Model Parameter ($\eta$) Total Instances Feasible & Bounded
Ratio (%)
Original Training Set (raw) N/A 1000 100.00%
ACM-MILP (acm-milp_ca) 0.05 1000 100.00%
0.1 1000 100.00%
0.2 1000 100.00%
DIG-MILP (fixed_digmilp_ca) 0.01 1000 100.00%
0.05 1000 100.00%
0.1 1000 100.00%
Table 12: Feasible and Bounded Ratio for Set Cover Instance Sets
Dataset / Model Parameter ($\eta$) Total Instances Feasible & Bounded
Ratio (%)
Original Training Set (raw/set_cover) N/A 1000 100.00%
G2MILP (g2milp_setcover) 0.01 1000 100.00%
0.05 1000 100.00%
0.1 1000 100.00%
Appendix G Branching nodes
Experiment Objective:
The goal is to quantify the difference in computational hardness between two sets of MILP instances: a ’training’ set and a ’generated’ set. This is achieved by measuring the total number of branch-and-bound (B&B) nodes explored by the Gurobi solver for each set and calculating the relative error between these totals.
Methodology & Operations:
  1. 1.
    Individual Instance Solving and Node Counting:
    • Gurobi Solver Configuration: For every instance processed, the solver operates under specific settings defined in the configuration:
      • Time Limit: A strict time limit of 300 seconds (solve.time_limit) is imposed on the solver for each instance. If the solver reaches this limit before completion, it stops, and the node count up to that point is used.
      • Threading: Each Gurobi solve process is restricted to using 1 CPU thread (solve.threads).
      • Solver Strategy Tuning: Additional parameters are set to potentially influence the B&B search and node count: MIPFocus is set to 1 (directing the solver to prioritize finding feasible solutions) and Heuristics is set to 0.5 (adjusting the effort spent on heuristic methods).
    • Node Count Retrieval: Upon completion or termination (e.g., hitting the time limit) of the Gurobi process for an instance, the total number of branch-and-bound nodes explored (NodeCount) is recorded. If an instance fails to process correctly or results in zero nodes (which is noted as potentially abnormal), its node count is typically excluded from the aggregation.
  2. 2.
    Total Node Aggregation:
    • The recorded node counts from all successfully processed instances in the training set are summed together to get a single value: $\sum\text{Nodes}_{\text{training}}$.
    • Likewise, the node counts from all successfully processed instances in the generated set are summed to get $\sum\text{Nodes}_{\text{generated}}$.
  3. 3.
    Relative Error Calculation:
    • The final metric, representing the relative difference in the total B&B effort between the two sets, is calculated using the specified formula:
      Relative $\text{Relative Error}=\left|\frac{\sum\text{Nodes}_{\text{generated}}-\sum\text{Nodes}_{\text{training}}}{\sum\text{Nodes}_{\text{training}}}\right|\times 100\%$
    • A check is performed to handle the case where $\sum\text{Nodes}_{\text{training}}$ is zero; in this scenario, the relative error cannot be meaningfully computed (often represented as infinity).
Key Parameters Influencing the Metric:
  • Gurobi Time Limit per Instance: 300 seconds.
  • Gurobi Threads per Instance: 1.
  • Gurobi MIPFocus Parameter: 1.
  • Gurobi Heuristics Parameter: 0.5.
Outcome:
This process yields a percentage value indicating how closely the total computational effort (measured in B&B nodes under specific solver settings and time limits) for the generated instances matches that of the training instances.
Table 13: Branching Node Statistics and Relative Error across Problem Types
Problem Model Mask Ratio ($\eta$) Mean Nodes Median Nodes Std Dev Max Nodes Time Limit Hits Relative Error (%)
IS Baseline (IS) N/A 14.3 1.0 49.1 1,056 0
G2MILP 0.01 26.2 1.0 60.1 724 0 84.1
G2MILP 0.05 765.7 175.0 1,941.6 16,314 5 5,271.6
G2MILP 0.10 8,386.4 8,330.0 1,219.5 11,789 982 58,730.0
ACM-MILP 0.05 8.8 1.0 36.8 572 0 38.3
ACM-MILP 0.10 10.9 1.0 63.8 1,344 0 23.4
ACM-MILP 0.20 12.5 1.0 78.0 1,977 0 12.3
CA Baseline (CA) N/A 18.5 1.0 48.1 514 0
ACM-MILP 0.05 6,556.0 1,661.5 9,926.3 51,913 138 35,360.0
ACM-MILP 0.10 6,504.9 1,547.0 9,962.8 54,385 120 35,084.3
ACM-MILP 0.20 4,873.6 962.5 8,021.7 43,036 75 26,259.8
DIG-MILP 0.01 34.1 1.0 76.2 799 0 84.4
DIG-MILP 0.05 34.3 1.0 76.4 812 0 85.6
DIG-MILP 0.10 37.8 1.0 83.5 902 0 104.3
SC Baseline (SC) N/A 46.4 1.0 298.6 8,135 0
G2MILP 0.01 85.4 1.0 536.5 14,592 0 84.0
G2MILP 0.05 79.6 1.0 430.9 10,643 0 71.6
G2MILP 0.10 63.7 1.0 269.8 4,713 0 37.3
Note: Statistics calculated over 1000 instances per model/ratio. Baselines refer to original training sets (Total Nodes: IS=14,255; CA=18,488; SC=46,408). ’Time Limit Hits’ indicates premature termination. Relative Error (RE) compares generated total nodes to baseline total nodes. $\eta$ denotes mask ratio.
Appendix H Structural Similarity
Experiment Objective:
The experiment aims to quantitatively compare the structural similarity between two distinct sets of MILP instances (referred to as Set 1 and Set 2). The comparison is based on extracting structural features from a graph representation of each instance and then evaluating the divergence between the feature distributions of the two sets using Jensen-Shannon Divergence (JSD).
Methodology & Operations:
  1. 1.
    Feature Extraction:
    • Graph Conversion & Feature Calculation (per instance): For each individual MILP instance file:
      • The instance is loaded and converted into a bipartite graph representation.
      • From this graph, a vector containing 11 predefined structural features is computed. These features are:
        1. (a)
          coef_dens: Coefficient density
        2. (b)
          var_degree_mean: Mean degree of variable nodes
        3. (c)
          var_degree_std: Standard deviation of variable node degrees
        4. (d)
          cons_degree_mean: Mean degree of constraint nodes
        5. (e)
          cons_degree_std: Standard deviation of constraint node degrees
        6. (f)
          lhs_mean: Mean of Left-Hand-Side coefficients (constraint matrix values)
        7. (g)
          lhs_std: Standard deviation of LHS coefficients
        8. (h)
          rhs_mean: Mean of Right-Hand-Side values (constraint bounds)
        9. (i)
          rhs_std: Standard deviation of RHS values
        10. (j)
          clustering: Graph clustering coefficient
        11. (k)
          modularity: Graph modularity
    • Result: This step produces two lists of feature vectors (NumPy arrays), one list corresponding to all instances in Set 1, and the other to all instances in Set 2.
  2. 2.
    Per-Feature Divergence & Similarity Calculation:
    • Before comparing the feature distributions, a sampling step is performed. However, in our experiment, the sampling number is set to 1000, which is also the total number of instances of both the original and generated sets.
    • The process iterates through each of the 11 feature dimensions (columns 0 through 10 in the sampled arrays f1 and f2).
    • For each feature $i$:
      • The script takes the 1000 values for this feature from Set 1 (f1[:, i]) and the 1000 sampled values from Set 2 (f2[:, i]).
      • The Jensen-Shannon Divergence (JSD) between these two empirical distributions is calculated (js_div function). This involves creating histograms (using 5 bins) for each set’s samples and for the combined samples, then applying the standard JSD formula based on Kullback-Leibler divergence.
      • The calculated JSD value ($JSD_{i}$) is transformed into a similarity score $S_{i}$ using the specified formula:
        $S_{i}=1-\frac{JSD_{i}}{\log(2)}$
      • This score $S_{i}$ ranges from 0 (maximal divergence) to 1 (identical distributions).
  3. 3.
    Overall Similarity Calculation:
    • After calculating the individual similarity scores ($S_{1},\ldots,S_{11}$) for all 11 features, the overall structural similarity score between Set 1 and Set 2 is computed.
    • This overall score is the simple arithmetic mean of the 11 individual feature similarity scores:
@@ Appendix H Structural Similarity
Key Computational Parameters:
  • Number of Parallel Workers for Feature Extraction: 64 (compute.num_workers)
  • Sample Size for JSD Calculation: 1000 (compute.num_samples) per dataset per feature.
Below should be the detailed results of all features.
Table 14: Overall structural similarity between generated Combinatorial Auction (CA) instances and the training dataset for different models and $\eta$ values.
Generative Model Masked Ratio ($\eta$) Overall Similarity
ACM-MILP 0.05 0.889
0.10 0.867
0.20 0.892
DIG-MILP 0.01 0.861
0.05 0.860
0.10 0.879
Table 15: Overall structural similarity between generated IS instances and the reference dataset for different models and masked ratios ($\eta$).
Generative Model Masked Ratio ($\eta$) Overall Similarity
ACM-MILP 0.05 0.734
0.10 0.731
0.20 0.699
G2MILP 0.01 0.466
0.05 0.473
0.10 0.486
Table 16: Overall structural similarity between generated Set Cover instances and the reference dataset for the g2milp model and different $\eta$ values.
Generative Model $\eta$ Value Overall Similarity
G2MILP 0.01 0.990
0.05 0.950
0.10 0.921
Appendix I Solving Time Gap

### v3

@@ Appendix A Datasets
A.1 Ecole Synthesized Instances
To facilitate various experiments, including model training and evaluation across different metrics, 4 synthetic datasets of MILP instances was generated. This process utilized the Ecole library (Prouvost et al., 2020), a platform designed for machine learning research in combinatorial optimization.
Parameter Randomization: Parameters like the number of constraints, variables, and density are randomly sampled for each instance within the ranges specified in the configuration file. This ensures diversity in the generated dataset. The details of synthesizing parameters are in Table 10.
@@ A.1 Ecole Synthesized Instances
A.2 Compatibility Testing with Public Datasets (NIPS ML4CO)
In addition to utilizing the synthetically generated dataset described above, the compatibility and robustness of the various analysis pipelines and tools developed (e.g., for feature extraction, solver log parsing, model evaluation) were verified using publicly available MILP benchmark instances. Specifically, datasets from the ML4CO Competition (Gasse et al., 2022) were employed for testing purposes. Successfully processing and running analyses on these standard benchmarks demonstrated the broader applicability and compatibility of the developed experimental framework beyond the specific synthetically generated instances. More results are in Appendix L.3
Appendix B More Related Work
@@
VAE-based Approaches: G2MILP (Geng and others, 2023) introduced the first DL-based framework, using a masked Variational Autoencoder (VAE) paradigm. It iteratively corrupts and replaces constraint nodes in the bipartite graph representation. While capable of generating instances structurally similar to the training data, G2MILP does not explicitly guarantee the feasibility or boundedness of the generated instances, potentially leading to unusable samples for certain downstream tasks. DIG-MILP (Wang and others, 2023), also VAE-based, addresses the feasibility and boundedness guarantee by leveraging MILP duality theories and sampling from a space including feasible solutions for both primal and dual formats. However, this can come at the cost of structural similarity compared to the original data. ACM-MILP (Guo and others, 2024) refines the VAE approach by incorporating adaptive constraint modification. It uses probability estimation in the latent space to select instance-specific constraints for modification, preserving core characteristics. It also groups strongly related constraints via community detection for collective modification, aiming to maintain constraint interrelations and improve hardness preservation.
Diffusion-based Approaches: Recognizing the power of diffusion models in generation tasks, MILP-FBGen (Zhang and others, 2024) proposed a diffusion-based framework. It uses a structure-preserving module to modify constraint/variable vectors and a feasibility/boundedness-constrained sampling module for the right-hand side and objective coefficients, aiming to ensure both structural similarity and feasibility/boundedness. Another diffusion-based approach focuses specifically on generating feasible solutions (rather than full instances) using guided diffusion. It employs contrastive learning to embed instances and solutions and uses a diffusion model conditioned on instance embeddings to generate solution embeddings, guided by constraints and objectives during sampling.
Structure-Focused Approaches: Some recent works explicitly target specific structural properties. MILP-StuDio (Liu and others, 2024) focuses on preserving and manipulating block structures commonly found in CCMs of real-world MILPs. It decomposes instances into block units, builds a library of these units, and uses operators (reduction, mix-up, expansion) to construct new, potentially scalable instances while maintaining structural integrity and properties like hardness and feasibility. MILPGen (Yang et al., 2024a)also aims at scalability, using node splitting/merging on the bipartite graph representation to decompose instances into tree structures, which are then concatenated to build larger problems.
@@ Appendix C Parameter Summary
See the parameters of our methods in Table LABEL:tab:exp_params_all.
Table 11: Summary of Key Experimental Parameters.
@@
Feasibility Ratio Gurobi time_limit 300 s Maximum Gurobi solve time per instance for feasibility/boundedness check.
Structural Similarity Feature Set 11 features Set of graph-based structural features extracted (e.g., density, degree stats).
Comparison Metric JSD-based Jensen-Shannon Divergence ($1-JSD/\log(2)$) averaged.
Branching Nodes Gurobi time_limit 300 s Maximum Gurobi solve time per instance.
Gurobi threads 1 Threads per solve (to enhance determinism).
Metric Relative Error $|(\sum N_{gen}-\sum N_{train})/\sum N_{train}|\times 100\%$.
Solving Time Gap Gurobi time_limit 3600 s Maximum Gurobi solve time per instance.
Gurobi threads 1 Threads per solve.
Metric Relative Difference $|(\text{mean\_t}_{gen}-\text{mean\_t}_{train})/\text{mean\_t}_{train}|\times 100\%$.
Hyperparameter Tuning Tuning Framework SMAC3 Automated algorithm configuration tool used.
n_trials 50 SMAC3 evaluation budget.
Objective Mean Solve Time Minimize mean wall-clock time on the tuning set.
Gurobi time_limit 60 s Gurobi time limit per solve during tuning/evaluation.
Gurobi threads 1 Gurobi threads per solve during tuning/evaluation.
Tuned Parameters Gurobi Controls 8 parameters including Heuristics, MIPFocus, etc.
Initial Basis Prediction GNN Model Parameters
@@
Gurobi threads 1 Gurobi threads per solve during evaluation.
@@
Solver-Internal Features Data Generation Phase
Gurobi time_limit 100 s Gurobi time limit for log generation.
@@
Root Gap Metric 1-Wasserstein Metric to compare scalar distributions.
Cut Plane Metric PCA + 1-W. Method for comparing high-dim cut vectors.
@@
PCA Scaling Standard Data scaling method (Z-score) before PCA.
Table 11: Summary of Key Experimental Parameters (Continued).
@@ Appendix E More Preliminaries
Simplex Method Basics
@@
The Simplex Method (Dantzig, 1963) is an algorithm primarily designed for solving Linear Programming (LP) problems, such as the LP relaxation defined above. While it can handle various forms, it often operates by converting the problem to a standard form like $\min\{\mathbf{c^{\prime}}^{\top}\mathbf{y}\mid\mathbf{A^{\prime}}\mathbf{y}=\mathbf{b^{\prime}},\mathbf{y}\geq 0\}$ (where $\mathbf{y}$ might include original and slack/surplus variables). It explores the vertices (Basic Feasible Solutions, BFS) of the feasible polyhedron $\mathcal{P}^{\prime}=\{\mathbf{y}\mid\mathbf{A^{\prime}}\mathbf{y}=\mathbf{b^{\prime}},\mathbf{y}\geq 0\}$. It iteratively moves between adjacent BFS by exchanging one variable in the current basis (a set of variables determining the BFS) with a non-basic variable, guided by criteria like reduced costs to ensure monotonic improvement of the objective function value, until an optimal BFS is found or unboundedness is detected.
Heuristics
@@
Cutting Planes (Cuts)
Cutting planes (Conforti et al., 2014) are linear inequalities, $\alpha^{T}\mathbf{x}\leq\beta$, that are valid for all feasible integer solutions, i.e., they hold for all $\mathbf{x}\in Conv(\mathcal{F}_{MILP})$ where $\mathcal{F}_{MILP}=\{\mathbf{x}\mid\mathbf{A}\mathbf{x}\leq\mathbf{b},x_{i}\in\mathbb{Z}\;\forall i\in\mathcal{I}\}$. However, they are chosen such that they are violated by the optimal solution $\mathbf{x}^{*}_{LP}$ of the current LP relaxation (i.e., $\alpha^{T}\mathbf{x}^{*}_{LP}>\beta$). Adding such cuts $\mathcal{C}$ to the LP relaxation feasible set $\mathcal{P}=\{\mathbf{x}\mid\mathbf{A}\mathbf{x}\leq\mathbf{b}\}$ yields a tighter relaxation $\mathcal{P}^{\prime}=\mathcal{P}\cap\{\mathbf{x}\mid\alpha^{T}\mathbf{x}\leq\beta$ for all, $\alpha^{T}\mathbf{x}\leq\beta\in\mathcal{C}\}$. The goal is to improve the lower bound derived from the relaxation, $\min_{\mathbf{x}\in\mathcal{P}^{\prime}}\mathbf{\tilde{c}}^{\top}\mathbf{x}\geq z_{LP}$, thereby bringing it closer to the true MILP optimum $z_{MILP}$ and potentially reducing the search space.
Appendix F Feasibility
Below is the result of feasibility test.
Table 12: Feasibility and Boundedness Ratios Across All Problem Types. Each configuration was tested on 1000 instances.
Problem Model Ratio ($\eta$) Feasible (%)
Independent Set Original Set N/A 100.00
G2MILP 0.01 100.00
0.05 100.00
0.10 93.40
ACM-MILP 0.05 100.00
0.10 100.00
0.20 100.00
Combinatorial Auction Original Set N/A 100.00
ACM-MILP 0.05 100.00
0.10 100.00
0.20 100.00
DIG-MILP 0.01 100.00
0.05 100.00
0.10 100.00
Set Cover Original Set N/A 100.00
G2MILP 0.01 100.00
0.05 100.00
0.10 100.00
Appendix G Structural Similarity
The experiment’s objective is to quantitatively compare the structural similarity between two sets of MILP instances. This is achieved by extracting graph-based structural features from each instance and then calculating the Jensen-Shannon Divergence (JSD) to measure the difference between the feature distributions of the two sets.
G.1 Methodology
1. Feature Extraction: Each MILP instance is converted into a bipartite graph. From this graph, a vector of structural features is extracted, capturing characteristics related to matrix sparsity, variable and constraint connectivity, coefficient statistics, and other graph-theoretic properties. This process is computationally parallelized for efficiency. The specific features are grouped into two categories and detailed in Table 13.
Table 13: Structural Features Used for Similarity Analysis
Category Feature Metric(s) Description
Basic Graph Features Coefficient Density coef_dens Reflects the sparsity of the constraint matrix.
Variable Node Degree var_degree_mean/std Describes the occurrence patterns of variables in constraints.
Constraint Node Degree cons_degree_mean/std Reflects the number of variables involved in each constraint.
Coefficient Statistics lhs/rhs_mean/std Captures the numerical distribution of constraint terms.
Advanced Topological Clustering Coefficient clustering Measures the local connection density between variables.
Modularity modularity Assesses the degree of modularity in the problem structure.
2. Per-Feature Similarity: For each feature, the Jensen-Shannon Divergence is used to compare its distribution across the two sets of instances. This divergence value is then normalized to a similarity score ranging from 0 (maximal divergence) to 1 (identical distributions).
3. Overall Similarity Score: A final, overall similarity score is calculated by taking the arithmetic mean of all the individual feature similarity scores. This single metric represents the aggregate structural similarity between the two instance sets.
Appendix H Solving Time Gap